Babygurl85:

When I was 6 my sister was half my age. If I'm 50 now then how old is my sister?

IrinaVal66:

easy 25

b1az34:

45

IrinaVal66:

because if u taake 50 divide it by 2 it equals 25.

Babygurl85:

@irinaval66 wrote:
easy 25
nope

b1az34:

42

Babygurl85:

@b1az34 wrote:
45
nope

IrinaVal66:

@b1az34 wrote:
42
no because 25+ 25 equals 50

Babygurl85:

@irinaval66 wrote:
@b1az34 wrote:
42
no because 25+ 25 equals 50
your still wrong

b1az34:

67

Babygurl85:

@b1az34 wrote:
67
no

IrinaVal66:

When u where 6 she was 3 so the number is cut right down the middle. 50-25 equals 25. So she is 25

Babygurl85:

@irinaval66 wrote:
When u where 6 she was 3 so the number is cut right down the middle. 50-25 equals 25. So she is 25
no

AlexisWolfe:

thats so easy its 56

Babygurl85:

@alexiswolfe wrote:
thats so easy its 56
no lol

hyuehswh:

easy 67 lmao you guys dont know math?

Babygurl85:

@hyuehswh wrote:
easy 67 lmao you guys dont know math?
ur wrong

hyuehswh:

nah im not but ok

Astrid1:

Your sister was half your age then. So you must've been 12 then- um hmm.

Astrid1:

GImme a second to piece it together.

IrinaVal66:

im not wrong its 25

b1az34:

@astrid1 wrote:
GImme a second to piece it together.
formula

IrinaVal66:

50 divided by 2

AlexisWolfe:

its super simple

Babygurl85:

is 47 bc she is only 3 years younger than me

hyuehswh:

but you said half ur age

IrinaVal66:

its

@hyuehswh wrote:
but you said half ur age
exactly

Rosalie:

47?

hyuehswh:

so you were 12

b1az34:

it was a riddel

hyuehswh:

not three years younger

Astrid1:

Wait you said half ur age so your sister was 3-

hyuehswh:

ohhh wait i read it wrong

Astrid1:

You were 6.

hyuehswh:

i just woke up

Astrid1:

So she's 47?

b1az34:

how?

Astrid1:

6/2 = 3 50-3 = 47

cutefacemya9:

omgggggg this is easy 20 becaus think about it she said HALF and half means two so basically its 20

b1az34:

so divide the nminus

Astrid1:

@astrid1 wrote:
So she's 47?
@babygurl85 ?

b1az34:

@astrid1 wrote:
6/2 = 3 50-3 = 47

b1az34:

@astrid1

Astrid1:

@b1az34 wrote:
@astrid1
WUT?

b1az34:

@b1az34 wrote:
@astrid1 wrote:
6/2 = 3 50-3 = 47
@astrid1

Babygurl85:

@hyuehswh wrote:
but you said half ur age
half of 6 is 3 and 3 is only 3 years younger than 6

Astrid1:

Why do you keep pinging me bro.

hyuehswh:

@babygurl85 wrote:
@hyuehswh wrote:
but you said half ur age
half of 6 is 3 and 3 is only 3 years younger than 6
k

Astrid1:

@babygurl85 wrote:
@hyuehswh wrote:
but you said half ur age
half of 6 is 3 and 3 is only 3 years younger than 6
SLAY I WAS RIGHT- SO SHES 47 RIGHT???

Babygurl85:

@astrid1 wrote:
@babygurl85 wrote:
@hyuehswh wrote:
but you said half ur age
half of 6 is 3 and 3 is only 3 years younger than 6
SLAY I WAS RIGHT- SO SHES 47 RIGHT???
yes

hyuehswh:

i Bananaing hate math

Babygurl85:

@hyuehswh wrote:
i Bananaing hate math
its not even math just common sense

hyuehswh:

i just Eviling woke up and that Evil was hella gooofy

Astrid1:

You just have to think outside of the box. Anyways, you should probably close this now that it's been solved.